OpenStudy (anonymous):

I need help, i have to graph this: cos^2x-2sinx*cosx-sin^2x=0 But i dont know how I would graph this. do I graph the factored form of this equation?

OpenStudy (anonymous):

use trig identities

OpenStudy (anonymous):

yea i got the factored form, i just want to know if i have to graph the factored form, cuz that way seems more easier

OpenStudy (anonymous):

what does it factor to?

OpenStudy (anonymous):

cuz i got (sinx+1)(2sinx-1)=0

OpenStudy (anonymous):

would i graph it like y=sinx +1 and then another graph y=2sinx-1?

OpenStudy (loser66):

Hello, Quebec friend

OpenStudy (anonymous):

lol hey ^_^

OpenStudy (anonymous):

im from Ontario :P

OpenStudy (anonymous):

i think you need to find a more simplified form. wolfram says this is equal to \[\sqrt{2} \sin \left( \frac{ \pi }{ 4 } - 2x\right)\]

OpenStudy (anonymous):

habs > your hockey teams ^_^ lol jk

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is that wat you got when u factored?

OpenStudy (anonymous):

i didn't really touch the problem. i typed the original equation in wolfram (a website) and got different forms of it. using trig id's you an apparently reduce it to that i can try figuring it out

OpenStudy (anonymous):

can*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill show u what i did

OpenStudy (anonymous):

\[\cos^{2}x-2sinx*cosx-\sin^{2}x=0\]\[1-\sin^{2}x-\sin2x-\sin ^{2}x=0\]\[-2\sin ^{2}x-\sin2x+1=0\]\[-1(\sin^{2}x-\sin2x+1)=0(-1)\]\[2\sin ^{2}x+\sin2x-1=0\]\[2\sin ^{2}x+2sinx-sinx-1=0\]\[(\sin^{2}x+2sinx)-(sinx-1)=0\]\[2sinx(sinx+1)-1(sinx+1)=0\] therefore \[(sinx+1)(2sinx-1)=0\] \[sinx+1=0\]\[sinx=-1\]\[x=\sin^{-1} (-1)\]\[x=\frac{ -\pi }{ 2 }\] or \[2sinx-1=0\]\[sinx=\frac{ 1 }{ 2 }\]\[x=\sin^{-1}( \frac{ 1 }{ 2 })\]\[x=\frac{ \pi }{ 6 }\]

OpenStudy (anonymous):

woo that took a while

OpenStudy (anonymous):

first i had to solve for x, and now i have to graph the equation

OpenStudy (anonymous):

i'm not sure that that is right. check this out: http://wolfr.am/1kEvDGL if you click "more" in the alternative forms place you'll see them i can't really help with this problem sorry

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

what do u think loser, do i graph the factored form? or the what euler said?

OpenStudy (loser66):

I don't know why you take a biiiiiiiiig U turn to back to original one like this, to me, cos^2 - 2 sinx cos x - sin^2 =0 rearrange cos^2 - sin^2 - 2 sinx cosx =0 and first 2 terms is cos (2x) so, it is cos (2x) - sin (2x) =0 and then solve for x if you want to get the roots, by the step cos (2x) = sin(2x) iff 2x = pi/4 + kpi If you want to graph, then, you don't have equal sign, I mean you have function f(x) = cos (2x) - sin (2x) from this, to graph, just give values to x and find value of f . for example x = 0, f (x) = 1 x = pi/2, f(x) = 0 x =pi , f(x) =.... to graph

OpenStudy (anonymous):

how did u get rid of 2sinx that was being multiplied by cosx?

OpenStudy (loser66):

2sinx cosx = sin 2x

OpenStudy (anonymous):

Yes, you have f(x), clumsy as it is. Choose some x values close together over the radian range 0 to 2 pi, evaluate f(x) there and plot them up, then join with smooth curve.

OpenStudy (anonymous):

yes and that wat i did

OpenStudy (loser66):

but your process is not correct at line 6, that step is NOT right

OpenStudy (anonymous):

to me that looks like a quadratic trigonometric equation

OpenStudy (loser66):

@douglaswinslowcooper hehehe, I am not useless as I am in physics, right?

OpenStudy (loser66):

@M0j0jojo not that, because sin 2x and 2 sinx is not the same

OpenStudy (anonymous):

ooh okay i see that now

OpenStudy (loser66):

for example, if x = pi/3 , then sin (2x) = sin (2pi/3) = sqrt(3)/2 and 2 sin x = 2* sqrt(3)/2 = sqrt(3)

OpenStudy (anonymous):

oh ok

OpenStudy (loser66):

good,

OpenStudy (anonymous):

i get confused between the two

OpenStudy (loser66):

not get, got

OpenStudy (loser66):

no more now

OpenStudy (anonymous):

yea lol thanks

OpenStudy (anonymous):

can i use the product and sum method on 2sin^2x+sin2x-1?

OpenStudy (anonymous):

i\[2\sin ^{2}x+\sin2x-1=0\] kind of lost my path of thought, so i how would go about solving for x from this point on

OpenStudy (anonymous):

u still there?

OpenStudy (anonymous):

this is a decent resource for trig id's: http://www.sosmath.com/trig/Trig5/trig5/trig5.html you could use the id: cos2x = 1 - 2sin^2(x) making what you have: sin2x - cos2x = 0 now solve for x